OpenStudy (anonymous):

Find the general formula for the derivative of function y = cos(2x)

OpenStudy (anonymous):

What I did 1. -2sin2x 2. -4cos2x 3.8six2x 4.16cos2x 5.-32cos2x 6.-64sin2x....etc so I found 2, -2, 2, -2, pattern.

OpenStudy (anonymous):

Using the chain rule: u(x) = 2x y(u) = cos(u) The chain rule is: dy/dx = (dy/du)(du/dx) dy/du = -sin(u) du/dx = 2 dy/dx = -sin(u) * 2 u = 2x dy/dx = -2sin(2x)

OpenStudy (anonymous):

I am looking for \[\frac{d^{n}}{dx^{n}} \]

OpenStudy (primeralph):